# From Knowledge Representation to Knowledge Organization and Back

### 12 Dec 2023 v1

@@ 3 Knowledge Organization (KO)
Thereafter, as illustrated in Figure 2, the canons of classification as briefed above are taken in input by the Classificationist to design (a set of) faceted classification schedules for either general knowledge organization usage, e.g., the Colon Classification ($C$) [44] designed and developed by Ranganathan himself (indicated with dashed lines), or, for specialized knowledge organization usage, e.g., Uniclass [^2] https://www.thenbs.com/our-tools/uniclass, specialized classification for the construction sector. Given the design of the classification schedule(s), the final activity of this phase shifts to the Classifier (see Figure 2) who has to now generate the unique Book Classification Number (Figure 2), e.g., Colon Number, for the subject matter of a book. To that end, the Classifier has to take in two important inputs. Firstly, (s)he has to take the a priori designed classification schedules which provide him/her with the concept hierarchy (with each concept uniquely identified via an identifier) as well as the formula (i.e., the facet formula) in which relevant concepts should be combined to generate the book classification number. Secondly, (s)he has to take in input the thought content of the book to be classified. Finally, the Classifier follows Ranganathan’s analytico-synthetic classification number generation procedure [44] to generate the Book Classification Number which uniquely identifies its subject matter. Notice that the above process is valid not only for books but also for any library resource.
@@ 4 From KR to KO and Back
Notice also, from Table 1, how the roles within the integrated KO-enriched KR methodology functionally compares with that of the original KR and facet-analytical KO methodology. The role of the Language Engineer (1.1), Knowledge Engineer (2.1) and User (3) in the integrated methodology is functionally similar to that of their counterparts in the KR and KO methodology (which have been functionally mapped before). The key difference, however, is reflected in the role of Ontology Engineer (1.2 + 2.3), which, builds the ETG hierarchy adhering to the canons (a completely new role absent in the other methodologies) and additionally specifies the data attributes within the ETG (part of the role of Data Scientist in the KR and the role of Catalog Code Designer in the KO methodology). The Data Scientist (2.2) in the KO-enriched KR methodology is functionally mapped to the role of Cataloguer in the KO and to a part of the role of Data Scientist in the KR methodology. This difference is because the former Data Sceintist (2.2) role only integrates the previous three outputs to generate the EG, whereas, the latter role of Data Sceintist (1.2 + 2.2), in addition to the above function, also specifies data attributes (performed by the new role of Ontology Engineer in the KO-enriched KR methodology). In terms of artifacts, the KO-enriched KR methodology is, functionally, the same as that of the KR methodology, and, thereof, to the KO methodology. Finally, in terms of activities, the KO-enriched KR methodology adds two extra activty flows over and above the KR methodology, namely: (i) the first activity flow from Ranganathan and Ongoing Work to the Ontology Engineer in terms of adapting guiding principles of $C$ and CCC, and (ii) the activity flow from the Ontology Engineer to the ETG Repository in terms of designing the repository.

### 09 Jan 2024 v2

@@ 3 Knowledge Organization (KO)
Thereafter, as illustrated in Figure 2, the canons of classification as briefed above are taken in input by the Classificationist to design (a set of) faceted classification schedules for either general knowledge organization usage, e.g., the Colon Classification (CC) [44] designed and developed by Ranganathan himself (indicated with dashed lines), or, for specialized knowledge organization usage, e.g., Uniclass [^2] https://www.thenbs.com/our-tools/uniclass, specialized classification for the construction sector. Given the design of the classification schedule(s), the final activity of this phase shifts to the Classifier (see Figure 2) who has to now generate the unique Book Classification Number (Figure 2), e.g., Colon Number, for the subject matter of a book. To that end, the Classifier has to take in two important inputs. Firstly, (s)he has to take the a priori designed classification schedules which provide him/her with the concept hierarchy (with each concept uniquely identified via an identifier) as well as the formula (i.e., the facet formula) in which relevant concepts should be combined to generate the book classification number. Secondly, (s)he has to take in input the thought content of the book to be classified. Finally, the Classifier follows Ranganathan’s analytico-synthetic classification number generation procedure [44] to generate the Book Classification Number which uniquely identifies its subject matter. Notice that the above process is valid not only for books but also for any library resource.
@@ 4 From KR to KO and Back
Notice also, from Table 1, how the roles within the integrated KO-enriched KR methodology functionally compares with that of the original KR and facet-analytical KO methodology. The role of the Language Engineer (1.1), Knowledge Engineer (2.1) and User (3) in the integrated methodology is functionally similar to that of their counterparts in the KR and KO methodology (which have been functionally mapped before). The key difference, however, is reflected in the role of Ontology Engineer (1.2 + 2.3), which, builds the ETG hierarchy adhering to the canons (a completely new role absent in the other methodologies) and additionally specifies the data attributes within the ETG (part of the role of Data Scientist in the KR and the role of Catalog Code Designer in the KO methodology). The Data Scientist (2.2) in the KO-enriched KR methodology is functionally mapped to the role of Cataloguer in the KO and to a part of the role of Data Scientist in the KR methodology. This difference is because the former Data Sceintist (2.2) role only integrates the previous three outputs to generate the EG, whereas, the latter role of Data Sceintist (1.2 + 2.2), in addition to the above function, also specifies data attributes (performed by the new role of Ontology Engineer in the KO-enriched KR methodology). In terms of artifacts, the KO-enriched KR methodology is, functionally, the same as that of the KR methodology, and, thereof, to the KO methodology. Finally, in terms of activities, the KO-enriched KR methodology adds two extra activty flows over and above the KR methodology, namely: (i) the first activity flow from Ranganathan and Ongoing Work to the Ontology Engineer in terms of adapting guiding principles of CC and CCC, and (ii) the activity flow from the Ontology Engineer to the ETG Repository in terms of designing the repository.